OpenStudy (emilyjones284):

MEDAL AND FAN! (: The equation x2 + 2y2 = 36 represents which conic section?

OpenStudy (emilyjones284):

hyperbola, circle, parabola, or ellipse?

OpenStudy (jdoe0001):

if the squared "x" and "y" are both positive, is an ellipse, or a circle, then again a circle is just a round ellipse

OpenStudy (emilyjones284):

the 2s are exponents...besides the one in front of the y

OpenStudy (jdoe0001):

well, when both are positive, doesn't matter who's in front of who it does only when one is negative, and not even then, because you could have -y + x ....

OpenStudy (anonymous):

That's a circle

OpenStudy (emilyjones284):

you guys sure its a circle?

OpenStudy (anonymous):

Yeap

OpenStudy (emilyjones284):

okay thanks (:

OpenStudy (jdoe0001):

hmmm

OpenStudy (jdoe0001):

hold the mayo

OpenStudy (anonymous):

I think it is elipse because it make much more sense

OpenStudy (jdoe0001):

is this what you have => x2 + 2y2 = 36? I mean or is it this => x2 + y2 = 36

OpenStudy (anonymous):

Hiper is x^2/a-y^2/b=1 Ellipse x^2/a+y^2/b=1 Circle x^2+y^2=r^2

OpenStudy (jdoe0001):

I just notice, the "y" has a coefficient of 2, not sure if it's a typo

OpenStudy (emilyjones284):

yes the coefficient of 2 is correct

OpenStudy (jdoe0001):

ahh, then is an ellipse

OpenStudy (jdoe0001):

hose the "2" and you get a circle hehe

OpenStudy (anonymous):

Example Graph the ellipse x2 y2 + = 1 25 16 Solution First identify a and b, remembering to take square roots. a = 5 b = 4 Next plot the vertices (-5,0), (5,0), (0,4), and (0,-4) Finally, connect the dots. The graph is shown to the right.

OpenStudy (anonymous):

Yea 2(x^2/2+y^2)=36 oops

OpenStudy (emilyjones284):

ohhh so its and ellipse not a circle?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

that's what I'm thinking

OpenStudy (anonymous):

Yea, I thought it was a typo too lol, srry

OpenStudy (anonymous):

wow that was close she almost got the wrong answer

OpenStudy (emilyjones284):

oh okay haha its okay thanks guys (: